OpenStudy (anonymous):

Two rooms in a house are similar rectangles. Room A is 12 ft by 16 ft. The longer side of Room B is 4 ft shorter than twice the length of the shorter side of Room A. What are the dimensions

OpenStudy (nincompoop):

longerside of B = 2(12)-4 B= 24-4 B = ?

OpenStudy (anonymous):

22

OpenStudy (anonymous):

20

OpenStudy (anonymous):

not 22

OpenStudy (nincompoop):

now that you have one of the dimensions of room B, we need to look for the other side. RECTANGLES have LENGTH and WIDTH correct? We can calculate the area of Room A to find out the other missing dimension of Room B, since they are similar rectangle - meaning either same area or same perimeter

OpenStudy (nincompoop):

how do we find the area of a rectangle @rach552 ?

OpenStudy (nincompoop):

@rach552 why did you close the question when it wasn't even finished yet?

OpenStudy (anonymous):

i fold it out